The problem is - A statue is 8ft tall. The display case for a model of statue is 18" tall. What scale allows for tallest model of statue that will fit in display case. I don't have any idea how to go about this problem.
Answer by edjones(7311)

8*12=96"
18/96=3/16 scale should do it.
.
Ed